A polling company reported that 57% of 1018 surveyed adults said that rising gas prices are quite annoying a. What is the exact value that is 57% of 1018? The exact value is Complete parts a through d below Type an integer or a decimal.) b. Could the result from part (a) be the actual number of adults who said that rising gas prices are quite annoying? Why or why not? O A. No, the result from part (a) could not be the actual number of adults who said that rising gas prices are quite annoying because that is a very rare opinion. B. Yes, the result from part (a) could be the actual number of adults who said that rising gas prices are quite annoying because the polling numbers are OC. No, the result from part (a) could not be the actual number of adults who said that rising gas prices are quite annoying because a count of people must result 。D. Yes, the result rom part a could be the actual number of adults who said hat rising gas prices are quite annoying because the results are statistically c. What could be the actual number of adults who said that rising gas prices are quite annoying? The actual number of adults with this opinion could be(Type an integer or a decimal.) d. Among the 1018 respondents, 418 said that rising gas prices are not at all annoying. What percentage of respondents said that rising gas prices are accurate. in a whole number. significant. not at all annoying? (Round to two decimal places as needed.)
Answer #1

a) Exact value = 0.57*1018 = 580.26

b) Option C is correct.

c) Actual number of adults = 580

d) Percentage = 418/1018 = 41.06 %
Answer #2
a)580.26. b) Yes . the results from part (a) could be number of adults who said that rising gas prices are “very harmful” beachside the results are statistically significant